# LaMSUM: Creating Extractive Summaries of User Generated Content using LLMs

### v2

@@ 3.4. Reimagining Summarization as an Election
Plurality voting and proportional voting are approval-based voting methods where voters can select multiple candidates they approve of without indicating a specific preference order. In multi-winner plurality voting (also known as block voting), each voter casts multiple votes and the candidates are selected based on the number of votes polled. In the context of summarization, a textual unit is treated as a candidate, and the LLM acts as the voter. We select the textual units in the decreasing order of the votes polled, till we obtain a summary of size $q$. PAV evaluates the satisfaction of each voter in the election outcome. A voter’s satisfaction is measured based on – amongst the number of candidates they voted for, how many are selected in the election. In the realm of summarization, PAV selects the textual units based on the amount of support each unit receives in $m$ summaries. Since both plurality and proportional are approval-based voting algorithms, the units are either approved or disapproved by the underlying LLM, with no explicit ranking or preference order. In this case, we prompt the LLM to select the best ¡$q$¿ sentences that summarize the input text as shown in Section 3.3.2 a.
@@
On the other hand, ranked choice voting entails assigning a score to each textual unit and subsequently selecting the highest-scoring units for inclusion in the summary. For ranked voting, we use the Borda count, a positional voting algorithm (Emerson, 2013). In the Borda method, each candidate is assigned points corresponding to the number of candidates ranked below them: the lowest-ranked candidate receives 0 points, the next lowest gets 1 point, and so forth. The candidates with the highest aggregate points are declared as the winners. In ranked voting, we prompt the LLM to output sentences in descending order of their suitability for the summary as discussed in Section 3.3.2 b.
@@
It is important to note that the prompting technique and the output generated by LLM vary for different voting methods. In approval voting the output from LLM is a list of $q$ textual units that LLM finds best suited to be included in the summary. Whereas in ranked choice voting, the output from LLM is a list of the same length as input i.e. $s$ with all the units sorted in decreasing order of their preference towards the summary, and Borda Count (Emerson, 2013) is used to identify the top $q$ textual units. In the next section, we highlight how the voting-based summarization schemes outperform the Vanilla setup, which does not use voting.
4. Experimental Setup
      Parameters Claritin US-Election MeToo
      #TextualUnits ($|\mathcal{T}|$) 3998 2107 483
#InputWords 53609 35522 16737
AV #CTU ($s$) 200 150 75
#CSTU ($q$) 100 100 50
RV #CTU ($s$) 40 40 40
#CSTU ($q$) 20 20 20
#SummaryTU ($k$) 100 100 50
@@ 4. Experimental Setup
Table 2. Input parameters used for the proposed framework LaMSUM. #TextualUnits is the number of textual units in the input set i.e. $|\mathcal{T}|$. #InputWords represents the number of words present in the input set $\mathcal{T}$. #CTU is the number of textual units in a chunk i.e. chunk size represented as $s$. #CSTU is the number of textual units in the chunk summary represented as $q$. #SummaryTU is the number of textual units present in the final summary i.e. $k$. AV and RV represents Approval Voting and Ranked Voting respectively.
4.1. Dataset
Our experiments are conducted on three publicly available datasets, consisting of crowd-sourced data from X, listed in Table 2 (Dash et al., 2019a). The Claritin dataset contains 3,998 tweets about the benefits and the side-effects of the anti-allergic drug Claritin. US-Election dataset contains 2,107 tweets from 2016 US Presidential Election where people support and attack different political parties. Me-Too dataset includes 483 tweets from the October 2018 MeToo movement, where individuals recount the harassment cases they experienced. [^3] Dataset: https://github.com/ad93/FairSumm/tree/master/Dataset During pre-processing, we remove all web links and duplicate entries from the datasets. The datasets also include gold-standard summaries (reference summaries) which are human-generated i.e., the textual units that are strong candidates for inclusion in the summary are selected by the humans. Claritin and US-Election dataset, each have three gold standard summaries, comprising 100 textual units each. Me-Too dataset has two gold-standard summaries with 50 textual units each.
4.2. Large Language Models (LLMs)
@@
LLMs are characterized by their extensive parameter sizes and remarkable learning abilities (Zhao et al., 2023; Chang et al., 2023). In our work, we utilize three open-source LLMs and one proprietary LLM to conduct experiments: llama3-70b-8192 from Meta (Touvron et al., 2023), mixtral-8x7b-32768 from Mistral AI (Jiang et al., 2024), gemini-1.0-pro from Google (Gemini Team, 2023) and
gpt-4o-mini-2024-07-18 from OpenAI (OpenAI, 2024). Across all experiments, we keep temperature, top probability and output tokens as 0, 0.8 and 8192 respectively.
4.3. Evaluation Metric
@@
For evaluating the quality of summaries generated by LaMSUM, we report ROUGE-1, ROUGE-2, and ROUGE-Lsum scores (Lin, 2004). ROUGE-1, ROUGE-2 and ROUGE-L respectively evaluate the overlap of unigrams, bigrams and longest common subsequence between the generated summary and the reference summary. ROUGE-Lsum is more suitable for extractive summarization, as it applies ROUGE-L at sentence level and then aggregates all the results to obtain the final score.
4.4. LaMSUM Input Parameters
@@
If $q\in[k,s)$, our proposed method can handle worst case scenario where all the textual units in the final summary may originate from a single chunk of level 0. As $q$ approaches $s$, more levels are required to converge to the final summary. The optimal value of $q$ that can handle worst case and also reduce the number of levels in multi-level summarization is $k$, thus we keep $q=k$ for experiments with approval voting. For instance, if $s$ is 200 and $q$ is 100, this indicates that only 50% of the units from each chunk advance to the next level.
@@ 5. Experimental Evaluation
                      Models Claritin US-Election         MeToo
R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum
                     ClusterRank 50.08 14.17 48.64 55.18 13.09 53.51 55.81 55.81 52.48
DSDR 28.01 8.40 27.45 48.04 9.53 46.44 55.03 17.30 52.04
LexRank 45.04 19.71 44.74 42.63 10.78 41.64 42.70 11.32 40.91
SummBasic 58.25 19.29 56.76 55.36 12.43 53.94 58.23 21.53 56.07
GPT2 61.61 23.58 60.74 55.86 15.07 54.81 40.63 11.24 38.86
BERT 57.30 22.37 56.21 55.89 15.44 55.00 45.72 10.76 43.50
XLNET 55.52 21.37 54.75 56.48 15.72 55.41 36.58 08.50 34.48
BERTSUM (Liu and Lapata, 2019) 57.87 22.75 55.96 59.00 17.51 57.41 57.11 23.08 54.84
 Vanilla Llama3 58.07 21.86 56.73 56.12 12.93 54.71 51.84 17.84 49.47
Mixtral 57.80 20.14 55.81 57.11 14.20 55.79 49.63 11.29 46.16
Gemini 49.70 18.33 48.78 51.00 12.36 49.19 45.51 8.95 42.51
GPT-4o mini 62.93 24.82 61.08 55.06 15.01 53.88 36.30 7.05 34.54
  Plurality Voting Llama3 61.28 23.79 59.54 60.11 18.26 58.99 55.51 18.79 52.85
Mixtral 59.13 22.90 57.40 59.55 16.01 58.12 55.41 17.62 52.63
Gemini 55.43 19.32 53.35 58.83 16.25 57.54 57.95 21.47 54.87
GPT-4o mini 64.20 26.71 62.66 58.15 15.91 56.78 54.55 19.66 51.9
Proportional Voting Llama3 61.31 23.36 59.73 58.90 15.99 57.55 58.14 14.99 55.46
Mixtral 60.30 24.03 58.58 58.46 15.74 57.28 57.36 22.97 54.82
Gemini 61.77 19.83 59.81 57.92 15.85 56.49 54.02 13.95 50.90
GPT-4o mini 64.13 26.75 62.30 58.68 19.08 57.62 48.55 13.23 45.35
 Borda Count Llama3 51.87 15.93 49.61 51.95 13.49 50.60 53.35 16.99 50.38
Mixtral 56.07 22.28 54.65 48.71 13.37 47.48 53.24 18.80 50.96
Gemini 51.45 16.92 49.39 54.01 13.38 52.56 50.13 21.45 47.96
GPT-4o mini 58.01 18.08 56.05 53.10 15.94 51.63 50.10 17.82 47.75
@@
Table 3. Table showing metric scores from different models for various datasets. Here, R1 = ROUGE-1 Score, R2 = ROUGE-2 Score, RLSum = ROUGE-LSum Score. The best value per dataset is shown in bold and clearly Approval voting outperforms all the other methods across all the evaluation measures.
5.1. Baseline Comparison
We compare LaMSUM with the pre-neural models (ClusterRank (Garg et al., 2009), DSDR (He et al., 2012), LexRank (Erkan and Radev, 2004), SummBasic (Nenkova and Vanderwende, 2005)), transformer based models (GPT2 (Radford et al., 2019), BERT (Miller, 2019), XLNET (Yang et al., 2019)) and with fine-tuned BERTSUM (Liu and Lapata, 2019) model. As presented in Table 3 it is observed that LaMSUM outperform the state-of-the-art summarization models. Earlier work (Zhang et al., 2023b) reported that the ChatGPT model achieves lower ROUGE scores on CNN/DM and XSum dataset. But our results demonstrate that our proposed framework LaMSUM performs significantly better than other fine-tuned models for large user-generated text.
Results from the LLM exhibit variability when executed multiple times over the same input. We conducted experiments with LLMs for five iterations for the same input. Table 3 contains the maximum ROUGE score obtained from these five iterations and Table 6 (in supplementary material) displays the variance in the ROUGE scores when executed for five times.
@@ 5.1. Baseline Comparison
     $R=\textsc{LLM}(\mathcal{T},si,ei,q)$triangleright $q$ textual units from $\mathcal{T}\in[si,ei]$
@@
5.2. Vanilla LLM vs. LaMSUM
Our proposed framework, LaMSUM, ensures robust summary generation by shuffling and employing a voting algorithm to select the best textual units for the summary. It is crucial to compare LaMSUM with a multi-level LLM that does not use shuffling and voting, which we call Vanilla LLM. Algorithm 3 outlines the steps used by vanilla LLM to find the chunk summary. Table 3 demonstrates that the vanilla multi-level LLM has lower ROUGE scores for each LLM compared to the proposed framework LaMSUM, indicating that shuffling and voting enhances the performance.
5.3. Which Voting Algorithm Holds the Lead?
We experimented with three voting algorithms, two approval-based and one ranked-based. Experimental results indicate that LLMs with approval voting perform the best compared to the ranked voting algorithm. We hypothesized that rank-based voting would yield better results, as it makes more informed decisions about the potential sentences to be included in the summary. Contrary to our expectations, rank-based algorithms performed even worse than neural and transformer-based models. This can be attributed to multiple factors: (i) LLMs hallucinate and output sentences in the same or in the reverse order as they were in the input. (ii) Occasionally, LLMs do not output all the sentences from the input, resulting in the padding of left-out sentences towards the end of the list, which disturbs the ranking and potentially affects the result. To overcome these problems, we kept the chunk size low as discussed in Section 4.4, but the results still did not surpass those of the approval-based voting algorithm.
Takeaway: LLMs, when prompted to select sentences that can summarize the input, perform better than when tasked to rank the sentences in the order of their preference towards the summary.
5.4. What Fails to Deliver Results?

### v4

@@ Introduction
Plurality voting and proportional voting are approval-based voting methods where voters can select multiple candidates they approve of without indicating a specific preference order. In multi-winner plurality voting (also known as block voting), each voter casts multiple votes and the candidates are selected based on the number of votes polled. In the context of summarization, a textual unit is treated as a candidate, and the LLM acts as the voter. We select the textual units in the decreasing order of the votes polled, till we obtain a summary of size $q$. PAV evaluates the satisfaction of each voter in the election outcome. A voter’s satisfaction is measured based on – amongst the number of candidates they voted for, how many are selected in the election. In the realm of summarization, PAV selects the textual units based on the amount of support each unit receives in $m$ summaries. Since both plurality and proportional are approval-based voting algorithms, the units are either approved or disapproved by the underlying LLM, with no explicit ranking or preference order. In this case, we prompt the LLM to select the best $<q>$ sentences that summarize the input text as shown in 1.
@@
On the other hand, ranked choice voting entails assigning a score to each textual unit and subsequently selecting the highest-scoring units for inclusion in the summary. For ranked voting, we use the Borda count, a positional voting algorithm (Emerson 2013). In the Borda method, each candidate is assigned points corresponding to the number of candidates ranked below them: the lowest-ranked candidate receives 0 points, the next lowest gets 1 point, and so forth. The candidates with the highest aggregate points are declared as the winners. In ranked voting, we prompt the LLM to output sentences in descending order of their suitability for the summary as discussed in 2.
@@
It is important to note that the prompting technique and the output generated by LLM vary for different voting methods. In approval voting the output from LLM is a list of $q$ textual units that LLM finds best suited to be included in the summary. Whereas in ranked choice voting, the output from LLM is a list of the same length as input i.e. $s$ with all the units sorted in decreasing order of their preference towards the summary, and Borda Count (Emerson 2013) is used to identify the top $q$ textual units. In the next section, we highlight how the voting-based summarization schemes outperform the Vanilla setup, which does not use voting.
Experimental Setup
Dataset
  No. Post Category A B C D E
 PC1 Rape/Sexual Assault 33 27 17 23 18
PC2 Chain Snatching/Robbery 49 16 103 32 22
PC3 Domestic Violence 27 87 30 10 34
PC4 Physical Assault 60 33 57 41 33
PC5 Stalking 146 166 165 100 153
PC6 Ogling/Staring 147 100 202 133 209
PC7 Taking Photos 73 70 43 37 55
PC8 Mas*****tion in public 45 55 27 30 42
PC9 Touching/Groping 152 209 206 159 230
PC10 Showing Po**ography 23 10 15 19 15
PC11 Commenting/Sexual Invites 133 192 273 133 131
PC12 Online Harassment 66 43 25 61 33
PC13 Human Trafficking 3 4 2 2 1
PC14 Others 50 23 36 41 16
Table 3: The distribution of posts across each category for five datasets – City A, City B, City C, City D, and City E. Posts on the Safe City platform are tagged with various categories by their authors, with each category representing a different form of sexual harassment.
                                                  Parameters City A City B City C City D City E
                                                  Total no. of textual units ($|\mathcal{T}|$) 625 867 866 545 728
Total no. of words in input set $\mathcal{T}$ 20544 12665 23501 22471 20807
Approval Voting No. of textual units in a chunk ($s$) 100 100 100 100 100
No. of textual units in chunk summary ($q$) 50 50 50 50 50
Ranked Voting No. of textual units in a chunk ($s$) 40 40 40 40 40
No. of textual units in chunk summary ($q$) 20 20 20 20 20
No. of textual units in final summary ($k$) 50 50 50 50 50
@@
Table 4: Input parameters used for the proposed framework LaMSUM. $|\mathcal{T}|$ is the number of textual units in the input set. $s$ is the chunk size i.e. the number of textual units in a chunk. $q$ is the number of textual units in the chunk summary. $k$ is the number of textual units present in the final summary.
We collect the Safe City posts from five different Indian cities [^4] The Red Dot Foundation has kindly made the data available for research purposes., which we refer to as City A, City B, City C, City D and City E. For City C and City E we obtain the posts for 3 years, i.e., from Dec 2021 to Nov 2024. For City A, City B and City D we consider the posts for 5 years, from Dec 2019 to Nov 2024. The rationale behind varying the duration of post selection is to keep the total number of posts below 1000, facilitating more efficient and accurate annotation by the human summarizers. Our dataset of City A, City B, City C, City D and City E consist of 625, 867, 866, 545 and 728 posts respectively. Sexual harassment can have various categories, such as physical assault, touching, stalking etc. Each post in the dataset is tagged with one or more categories by the author of the post. The distribution of posts across these categories is shown in Table 3. While the Safe City platform does not collect personal information such as names or identities, it does gather age, gender, and details of the incidents. Table 8 (in the Appendix) provides an overview of the attributes associated with each post. For our task, we focus solely on the main description provided in the posts.
For each of the five city-specific datasets, we generate gold-standard (reference) summaries created by three domain experts. These experts carefully selected textual units from the posts that are strong candidates for inclusion in the summary. As a result, each dataset has three expert-generated gold-standard summaries, each comprising 50 textual units (posts). The expert annotators were provided with clear guidelines for selecting the posts for the reference summary – i) Diversity: Prioritize diverse posts that represent various forms of assault. ii) Descriptive: Give preference to posts with detailed descriptions over those containing only 2-3 words. iii) Severity: Include posts that depict more serious cases or require urgent attention. iv) Redundancy: Exclude posts that are repetitive or redundant. Additional details about the posts selected by the annotators is available in Figure 7 and Table 10 in the Appendix .
Large Language Models (LLMs)
@@
LLMs are characterized by their extensive parameter sizes and remarkable learning abilities (Zhao et al. 2023; Chang et al. 2023). In our work, we utilize two open-source LLMs and one proprietary LLM to conduct experiments: llama-3.1-8B-instruct from Meta (Touvron and et al. 2023), open-mistral-nemo-2407 from Mistral AI (Jiang and et al. 2024) and gpt-4o-mini-2024-07-18 from OpenAI (OpenAI 2024). Across all experiments, we keep temperature, top probability and output tokens as 0, 0.8 and 8192 respectively. We downloaded llama-3.1-8B-instruct from Hugging Face and executed it locally on A6000 GPU. For open-mistral-nemo and gpt-4o-mini, we utilised official APIs.
Evaluation Metric
@@
For evaluating the quality of summaries generated by LaMSUM, we report ROUGE-1, ROUGE-2, ROUGE-Lsum (Lin 2004). ROUGE-1, ROUGE-2 and ROUGE-L respectively evaluate the overlap of unigrams, bigrams and longest common subsequence between the generated summary and the reference summary. ROUGE-Lsum is more suitable for extractive summarization, as it applies ROUGE-L at sentence level and then aggregates all the results to obtain the final score.
LaMSUM Input Parameters
@@
If $q\in[k,s)$, our proposed method can handle worst case scenario where all the textual units in the final summary may originate from a single chunk of level 0. As $q$ approaches $s$, more levels are required to converge to the final summary. The optimal value of $q$ that can handle worst case and also reduce the number of levels in multi-level summarization is $k$, thus we keep $q=k$ for experiments with approval voting. For instance, if $s$ is 100 and $q$ is 50, this indicates that only 50% of the units from each chunk advance to the next level.
@@
  City A City B City C City D City E
Models R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum
 LexRank 57.644 28.084 55.904 38.898 16.909 38.239 56.379 24.897 54.057 46.221 21.709 44.612 52.157 23.681 49.812
SummBasic 52.498 21.848 50.700 26.714 9.536 26.283 53.172 21.600 51.136 43.215 19.841 41.392 50.677 21.357 48.482
LSA 55.250 25.156 53.219 52.529 25.947 51.462 45.623 17.378 43.981 42.370 18.262 40.762 47.522 17.521 45.585
BERT 55.645 23.673 54.236 53.799 26.428 52.711 53.610 22.353 51.982 43.986 19.738 42.322 50.444 21.470 49.377
XLNET 50.559 19.373 48.959 51.706 24.259 50.937 52.819 21.137 51.189 43.768 22.133 42.391 51.650 22.151 50.142
BERTSUM 56.570 25.323 54.226 51.137 24.551 49.386 53.927 22.647 50.993 53.767 25.846 49.376 47.735 20.606 45.295
LaMSUM 60.440 30.006 58.355 56.296 32.109 55.229 58.708 29.022 56.428 54.855 28.124 52.504 55.921 29.919 53.426
@@
Table 5: Table showing metric scores from different models for various datasets. Here, R1 = ROUGE-1 Score, R2 = ROUGE-2 Score, RLSum = ROUGE-LSum Score. The best result for each dataset is shown in bold and clearly LaMSUM outperforms all the other methods across all the evaluation measures.
Baseline Comparison
We compare LaMSUM with the pre-neural models such as (LexRank (Erkan and Radev 2004), SummBasic (Nenkova and Vanderwende 2005), LSA (Gong and Liu 2001)), transformer based models (BERT (Miller 2019), XLNET (Yang et al. 2019)) and with state-of-the-art fine-tuned BERTSUM (Liu and Lapata 2019) model. Our proposed method, LaMSUM, achieves optimal performance when gpt-4o-mini is used with Proportional Approval Voting. Therefore, we present the results of this combination as the outcomes of LaMSUM. As shown in Table 5, it is observed that LaMSUM surpasses state-of-the-art summarization models across all metrics.
Does LaMSUM perform better than Vanilla LLM?
Figure 4: Metric scores obtained through four different LLM setups. (i) Vanilla LLM without shuffling and voting method (ii) LaMSUM with Plurality Approval Voting (iii) LaMSUM with Proportional Approval Voting and (iv) LaMSUM with Borda Count Ranked Voting. Results demonstrate that gpt-4o-mini with Proportional Approval Voting performs the best across all the cases. Here, Llama, Mixtral and GPT refers to llama-3.1-8B, open-mistral-nemo, and gpt-4o-mini respectively.
Our proposed framework, LaMSUM, ensures robust summary generation by shuffling and employing a voting algorithm to select the best textual units for the summary. It is crucial to compare LaMSUM with a multi-level LLM that does not use shuffling and voting, which we call Vanilla LLM. Algorithm 3 outlines the steps used by vanilla LLM to find the chunk summary. Figure 4 demonstrates that the vanilla multi-level LLM has lower ROUGE scores for each LLM compared to the proposed framework LaMSUM, indicating that shuffling and voting enhances the performance. Earlier work (Zhang, Liu, and Zhang 2023b) reported that the ChatGPT model achieves lower ROUGE scores on CNN/DM and XSum dataset. But our results demonstrate that our proposed framework performs significantly better than other fine-tuned language models such as BERTSUM for large user-generated text.
Which LLM takes the lead?
We conducted experiments using three LLMs – llama-3.1-8B, open-mistral-nemo, and gpt-4o-mini. As shown in Figure 4, gpt-4o-mini demonstrated the best performance among the three. While we utilized smaller models for these experiments, leveraging larger LLMs with higher parameter counts is likely to result in even better performance.
Which voting algorithm performs the best?
We experimented with three voting algorithms, two approval-based (pluarlity and proportional) and one ranked-based (borda-count). Experimental results indicate that LLMs with proportional approval voting perform the best compared to the other voting algorithms. As shown in Figure 4, it can be observed that PAV performs the best across all the LLMs. We hypothesized that rank-based voting would yield better results, as it makes more informed decisions about the potential sentences to be included in the summary. Contrary to our expectations, rank-based algorithms did not surpass the proportional approval voting. This can be attributed to multiple factors: (i) LLMs may hallucinate and output sentences in the same or in the reverse order as they were in the input. (ii) Occasionally, LLMs do not output all the sentences from the input, resulting in the padding of left-out sentences towards the end of the list, which disturbs the ranking and potentially affects the result. To overcome these problems, we kept the chunk size low, but the results still did not surpass those of the proportional approval-based voting algorithm.
@@
     $R=\textsc{LLM}(\mathcal{T},si,ei,q)$triangleright $q$ textual units from $[si,ei]$
@@
Analyzing the LaMSUM output
Figure 5: Posts chosen by LaMSUM tend to be detailed and descriptive, offering a deeper level of information. Number of words in LaMSUM selected posts is often highest across various datasets, ensuring extensive and comprehensive summarization.
  Models City A City B City C City D City E
 LexRank 7.486 5.819 7.637 7.548 7.481
SummBasic 8.198 5.734 8.050 8.020 8.196
LSA 8.251 7.061 8.481 8.194 8.387
BERT 8.068 7.762 8.437 8.186 8.191
XLNET 8.144 7.689 8.619 8.205 8.400
BERTSUM 8.331 7.539 8.690 7.960 8.112
LaMSUM 8.563 7.822 8.622 8.606 8.480
Table 6: Entropy values representing the diversity in the summaries produced by various algorithms. Bold values highlight the best and the underline denotes the second-best performance. LaMSUM achieves the highest diversity score across four datasets.
We analyse the difference between the posts selected by the LaMSUM and those chosen by other algorithms. LaMSUM selects the post that are more descriptive and rich in detail, in contrast to posts that lack sufficient information. As shown in Figure 5, it is evident that across nearly all datasets, the posts selected by LaMSUM exhibit a higher word count. This indicates that the proposed algorithm is capable of capturing more detailed information compared to the others algorithms (refer Table 11 in Appendix for examples). Furthermore, posts selected by LaMSUM exhibit greater diversity, encompassing a broader range of harassment categories compared to the other baselines. We use entropy as a measure of diversity where a higher value is indicative of more randomness (Jost 2006). Table 6 demonstrates that LaMSUM generates diverse summaries as compared to other baselines.
Accompanying Website
We have demonstrated so far that LaMSUM generates superior summaries for Safe City posts. Building on this, we apply LaMSUM to posts from various localities and build a companion website for Safe City. This website enables the end-users and authorities to quickly access an overview of incidents in a specific area and customized time frame, enhancing awareness and helping policy action. For simplicity, we segregate the posts selected in the summary on the basis of the post categories. Figure 6 presents a screenshot of our accompanying website, which aims to assist users in making more informed decisions. [^5] The website link is hidden to maintain anonymity.
Figure 6: Homepage of the accompanying website which allows users to obtain a quick snapshot of the incidents happening in an area. Results can be customized as per the district and the timeframes.
Concluding Discussion
Safe City platform receives numerous posts related to sexual harassment, having a summarization algorithm enables end-users to quickly review significant posts. This work marks an early attempt to achieve extractive summarization of large user-generated text that exceeds a single context window using zero-shot learning. The proposed multi-level framework LaMSUM leverages approval based and ranked based voting algorithms to generate robust summaries. Experiments conducted on Safe City crowd-sourced dataset demonstrated the efficacy of LaMSUM, as it outperformed the results achieved by state-of-the-art models.
Note that there can be a concern regarding the potential data leakage, as the experiments involve newer LLMs that may have been exposed to the experimented datasets during their pre-training phase. We showcased that the vanilla LLM, which also includes LLM underperformed, whereas our proposed framework which generates robust summaries yielded good results. This highlights the efficacy of our model, even when it is exposed to data leakage.
Limitation. Our proposed framework, LaMSUM, very well handles text of any length, conditioned on the fact that the final summary fits within a single context window. Some modifications to LaMSUM may be necessary when the output summary exceeds the size of a single context window.
Ethical Considerations. Our research focuses on using LLMs to produce extractive summaries for platforms like Safe City. LLMs often exhibit bias towards their training data, which can influence their preference for certain textual elements during the summarization process. Their “black box” nature, with an opaque decision-making process, makes it challenging to discern how or why specific textual units are chosen for summarization. The posts selected by LaMSUM may not accurately represent real-life scenarios and cannot serve as a reliable proxy for actual situations. Additionally, LaMSUM may overlook less frequent posts with limited informational content. As a result, exclusive reliance on our framework could lead to the oversight of specific issues by the authorities. While LaMSUM can produce high-quality summaries, its use must be approached with careful consideration of potential ethical implications.
Appendix A Appendix
Algorithm
Algorithm 4 shows the method for output calibration by utilizing two modules – i). minimum edit distance and ii). maximum count of keywords.
Algorithm 4 Algorithm for output calibration
function Check($R,\mathcal{T},si,ei$)
     $Y=\{\}$triangleright store the result
     for $x$ in $R$ do $\triangleright$ for each sentence $x$ in LLM result $R$
         $min\_dist=\infty$triangleright keep track of min distance
         $min\_idx=-1$triangleright post with min edit distance
         $max\_count=0$triangleright matching keywords
         $max\_idx=-1$triangleright post with max keywords
         $K=\textsc{Keywords}(x)$triangleright obtain keywords in $x$
         for $i\leftarrow si$ to $ei$ do $\triangleright$ for each unit in $\mathcal{T}$
              $d=\textsc{EditDist}(x,\mathcal{T}[i])$triangleright obtain edit distance
              if $d<min\_dist$ then $\triangleright$ lesser edit distance
                  $min\_dist=d$triangleright update $min\_dist$
                  $min\_idx=i$triangleright update $min\_idx$
              end if
              $c=\textsc{Count}(K,\mathcal{T}[i])$triangleright # keywords in $\mathcal{T}[i]$
              if $c>max\_count$ then $\triangleright$ lesser edit distance
                  $max\_count=c$triangleright update $max\_count$
                  $max\_idx=i$triangleright update $max\_idx$
              end if
         end for
         if $min\_dist<\epsilon$ then $\triangleright$ edit distance is low
              $Y.add(\mathcal{T}[min\_idx])$triangleright add to result $Y$
         else
              $Y.add(\mathcal{T}[max\_idx])$triangleright add to result $Y$
         end if
     end for
     return $Y$
end function
Zero Shot Prompting
  Prompts
 1 Select the most suitable units that summarize the input text.
Prompt: Input consists of $<$chunk_size$>$ sentences. Each sentence is present in a new line. Each sentence contains a sentence number followed by text. You are an assistant that selects best $<$summary_length$>$ sentences (subset) which summarizes the input. Think step by step and follow the instructions. $<$sentences$>$
 2 Generate a ranked list in descending order of preference.
Prompt: Input consists of $<$chunk_size$>$ sentences. Each sentence is present in a new line. Each sentence contains a sentence number followed by text. You are an assistant that selects best $<$summary_length$>$ sentences (subset) which summarizes the input. Think step by step and follow the instructions. $<$sentences$>$
Table 7: Prompts utilised for 1 Approval and 2 Ranked based voting algorithm.
What Fails to Deliver Extractive Summary ?